OpenStudy (anonymous):

Find the exact real value of arccos (-√2/2)

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

pi divided by four three pi divided by four seven pi divided by three negative pi divided by four

OpenStudy (jdoe0001):

\(\bf cos^{-1}\left(-\frac{\sqrt{2}}{2}\right)\ \quad \\\implies cos(\theta)=-\cfrac{\sqrt{2}}{2}\implies cos^{-1}(cos(\theta))=cos^{-1}\left(-\frac{\sqrt{2}}{2}\right)\\ \quad \\ \theta=cos^{-1}\left(-\frac{\sqrt{2}}{2}\right)\) check your Unit Circle

OpenStudy (anonymous):

3pi/4

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

Thank you(((: